Solve for the variable: 2x2^ + 8x = 0
a) 0 or 4
b) 0 only
c) 4 only
d) 0 or -4
e) None of these

Answer by checkley71(8403)

2X^2+8X=0
2X(X+4)=0
2X=0
X=0 SOLUTION
X+4=0
X=-4 SOLUTION
SOLUTIONS ARE (0 & -4)